OpenStudy (twizttiez):

Solve the given system, using the substitution method. y = 3x – 4 9x – 3y = 14 A. There is no solution. B. (12, 14) C. (14, 12) D. There are an infinite number of solutions.

OpenStudy (twizttiez):

@DanJS

OpenStudy (twizttiez):

i think its a

OpenStudy (solomonzelman):

yes that is correct, no solution:

OpenStudy (solomonzelman):

\(\large\color{black}{ y = 3x - 4 \\ 9x -3y = 14 }\) \(\large\color{black}{ y = 3x - 4 \\ -3y = -9x+14 }\) \(\large\color{black}{ y = 3x - 4 \\ y = 3x-\frac{14}{3} }\)

OpenStudy (solomonzelman):

they are parallel lines that never intersect.... good work.

OpenStudy (danjs):

if the x and y terms are a scalar multiple of each other, the lines are parallel, if the term on the other side of the = sine is also the same multiple, they are the same line.